# Dynamical Similarity in Multisymplectic Field Theory

### v3

@@ IV.1 Contact Reduction of Lagrangian Systems
In this construction, we have examined only the equation of motion for $\rho$; however, if we wish to affirm that our symmetry-reduced theory is dynamically-equivalent to that described by the original multisymplectic Lagrangian $\widehat{L}$, we must also verify that the equations of motion of the unscaled variables are faithfully reproduced. In order to do so, we should take into account the fact that, within the symmetry-reduced space, the velocity coordinate $\rho_{\mu}$ is to be replaced with the appropriate expression, in terms of $\psi$, $\psi_{\mu}$, and $s^{\mu}$ [25]. Thus, in the Herglotz-Lagrange equation for $\psi$
$\frac{\partial}{\partial x^{\mu}}\frac{\partial L^{H}}{\partial\psi_{\mu}}-\frac{\partial L^{H}}{\partial\psi}-\frac{\partial L^{H}}{\partial\psi_{\mu}}\frac{\partial L^{H}}{\partial s^{\mu}}=0$ (54)
we have the following terms
$\begin{split}\frac{\partial L^{H}}{\partial\psi_{\mu}}=\frac{\partial f}{\partial\psi_{\mu}}+\frac{\partial f}{\partial\rho_{\nu}}\frac{\partial\rho_{\nu}}{\partial\psi_{\mu}}-s^{\nu}\frac{\partial\rho_{\nu}}{\partial\psi_{\mu}}\\ \frac{\partial L^{H}}{\partial\psi}=\frac{\partial f}{\partial\psi}+\frac{\partial f}{\partial\rho_{\nu}}\frac{\partial\rho_{\nu}}{\partial\psi}-s^{\nu}\frac{\partial\rho_{\nu}}{\partial\psi}\\ \frac{\partial L^{H}}{\partial s^{\mu}}=\frac{\partial f}{\partial\rho_{\nu}}\frac{\partial\rho_{\nu}}{\partial s^{\mu}}-\rho_{\mu}-s^{\nu}\frac{\partial\rho_{\nu}}{\partial s^{\mu}}\end{split}$
@@
Combining these results, and recalling from (53) that $s^{\mu}=\partial f/\partial\rho_{\mu}$, we find that the left-hand side of (54) becomes
$\frac{\partial}{\partial x^{\mu}}\frac{\partial f}{\partial\psi_{\mu}}-\frac{\partial f}{\partial\psi}+\rho_{\mu}\frac{\partial f}{\partial\psi_{\mu}}$
Upon multiplying this expression by $e^{\rho}$, it is apparent that the Herglotz-Lagrange equation (54) implies that
$\frac{\partial}{\partial x^{\mu}}\left(e^{\rho}\frac{\partial f}{\partial\psi_{\mu}}\right)-e^{\rho}\frac{\partial f}{\partial\psi}=0$
which is precisely the multisymplectic field equation for $\psi$. Consequently, we have demonstrated that the Herglotz Lagrangian calculated according to (53) correctly reproduces the dynamics of the original system; moreover, this new description makes no reference to $\rho$, confirming our statement that this scaling variable plays no role in the dynamical evolution of the physically-observable degrees of freedom. The symmetry-reduced system is then multicontact or pre-multicontact, depending upon the Lagrangian form $\Theta_{L^{H}}$; a detailed discussion of the geometry of the symmetry-reduced space is postponed until section (IV.2), having presented both the Lagrangian and Hamiltonian formalisms.
We conclude this section with a number of remarks: firstly, the procedure we have outlined above is often far more straightforward in practice. The vast majority of field theories of interest, which admit dynamical similarities, such as scalar field theories, bosonic string theory, and even General Relativity, do not require spacetime coordinate reparameterisations; in this case, we modify the reduction process slightly. The vector fields (45) and 46) remain valid, setting $A=0$; we also seek a scaling function $\xi=\phi^{1/B}$, as before; however, instead of identifying $\xi=e^{\rho}$, we must take $\xi=e^{\rho/\Lambda}$. This serves to ensure that the Lagrangian acquires the final form $L=e^{\rho}f(\rho_{\mu},\psi,\psi_{\mu})$, which is essential for the subsequently-derived Herglotz Lagrangian to produce the correct on-shell dynamics.
Finally, we must also address the case of degenerate Lagrangians [42, 81]; one of our principal objectives in developing this mathematical framework is the study of classical General Relativity. It has been demonstrated that, upon making a conformal decomposition of the spacetime metric, the Einstein-Hilbert Lagrangian admits a scaling symmetry, in which the scaling variable is precisely the conformal factor [25]. This analysis was performed within the second-order formalism, framed in the language of $k$-symplectic geometry [82]. It will be of great interest, therefore, to examine the details of this dynamical similarity in the first-order vielbein formalism [83, 84]. In addition to General Relativity, any other gauge theory we may wish to study will also require the development of suitable geometric tools, adapted to the treatment of singular Lagrangians [85]. The symmetry reduction of such theories is an area of active research; it is, at present, not clear, for example, if the process of gauge-fixing commutes with the elimination of the scaling variable. Additionally, there remain several open questions regarding the topology of the reduced space, when the original pre-multisymplectic theory is confined to a constraint submanifold of the covariant phase space.
IV.2 Contact Reduction of Hamiltonian Systems
In the previous section, we presented an algorithmic procedure for the symmetry reduction of Lagrangian systems admitting a dynamical similarity; we shall now provide the complementary Hamiltonian construction, together with a detailed discussion of the geometry of the reduced space, within both the Lagrangian and Hamiltonian formalisms. The reduction procedure for the standard symplectic case has been formulated rigorously in [72], and our objective is to generalise these principles to a field-theoretic context.
Suppose that $(J^{1}E^{*},\Omega_{h})$ is a multisymplectic Hamiltonian system (over the $d$-dimensional spacetime manifold $M$), with corresponding Hamiltonian function $H:J^{1}E^{*}\rightarrow\mathbb{R}$. For simplicity, we shall also temporarily assume that the Legendre map is a global diffeomorphism, so that our system is hyperregular. Recall, from section (II.3), that we constructed the 2-forms [^6] Here, we have adapted our notation, so as to reflect the fact that spacetime indices assume values $\mu=0,\,\cdots,d-1$.
$\omega^{\mu}_{h}=-\,\iota_{\partial_{d-1}}\cdots\;\iota_{\partial_{0}}(\Omega_{h}\wedge dx^{\mu})$ (55)
We shall declare a vector field $D\in\mathfrak{X}^{\infty}(J^{1}E^{*})$ to constitute a dynamical similarity of the Hamiltonian system if
$\mathfrak{L}_{D}H=\Gamma H\quad\quad\quad\textrm{and}\quad\quad\quad\mathfrak{L}_{D}\omega^{\mu}_{h}=\omega_{h}^{\mu}\quad\textrm{for all}\;\mu=0,\cdots\,,d-1$
As in the Lagrangian case, we refer to $\Gamma\in\mathbb{R}$ as the degree of the symmetry, and employ the notation $D$ and $\Gamma$, so as to make a distinction from the corresponding Lagrangian objects $\Sigma$ and $\Lambda$.
@@ IV.2 Contact Reduction of Hamiltonian Systems
The Hamiltonian reduction may, in principle, be carried out without reference to the Lagrangian function; however, in practice, it is more convenient to work with the Hamiltonian obtained as a result of a Legendre transform of $\widehat{L}$, introduced in (52). Writing
$\widehat{L}=e^{\rho}f(\rho_{\mu},\psi,\psi_{\mu})$
we see that
$\mathcal{FL}^{*}p^{\mu}_{\rho}=\frac{\partial\widehat{L}}{\partial\rho_{\mu}}=e^{\rho}\frac{\partial f}{\partial\rho_{\mu}}=e^{\rho}s^{\mu}\quad\quad\quad\quad\mathcal{FL}^{*}p^{\mu}_{\psi}=\frac{\partial\widehat{L}}{\partial\psi_{\mu}}=e^{\rho}\frac{\partial f}{\partial\psi_{\mu}}$ (56)
@@
Recall that in the coordinates $(\rho,\rho_{\mu},\psi,\psi_{\mu})$, the Lagrangian scaling symmetry adopted the particularly simple form $\Sigma=\partial_{\rho}$; by virtue of this structure, we find that the corresponding vector field $D$ (with $\Gamma=1$) for the Hamiltonian is
$D=\frac{\partial}{\partial\rho}+p^{\mu}_{\rho}\frac{\partial}{\partial p^{\mu}_{\rho}}+p^{\mu}_{\psi}\frac{\partial}{\partial p^{\mu}_{\psi}}$ (57)
For the case of multiple unscaled fields $\psi^{a}$ ($a=1,\,\cdots,k$), this vector field generalises trivially as the sum of logarithmic derivatives of the momenta.
In order to obtain the symmetry-reduced Hamiltonian, we introduce (omitting the pullback by the appropriate submersion)
$s^{\mu}=\frac{p^{\mu}_{\rho}}{e^{\rho}}\quad\quad\quad\Pi^{\mu}_{\psi}=\frac{p^{\mu}_{\psi}}{e^{\rho}}$ (58)
Finally, we claim that, upon defining [^7] Here, we have made explicit the dependence of $H^{c}$ and $\eta$ on $\psi$, $\Pi^{\mu}_{\psi}$, and $s^{\mu}$ to emphasise that these are to be considered the coordinates on the reduced space.
$H^{c}(\psi,\Pi^{\mu}_{\psi},s^{\mu})=\frac{H}{e^{\rho}}\quad\quad\quad\quad\eta^{\mu}(\psi,\Pi^{\nu}_{\psi},s^{\nu})=\frac{\iota_{D}\omega^{\mu}_{h}}{e^{\rho}}=ds^{\mu}-\Pi^{\mu}_{\psi}\,d\psi$ (59)
@@
we obtain a (pre-)multicontact structure, whose corresponding Hamiltonian $d$-form $\Theta_{H^{c}}$ is given by
@@
Moreover, as we shall show, the Herglotz De-Donder Weyl equations for holonomic sections correctly reproduce the dynamics derived from the full multisymplectic system, and further, they do so without reference to the scaling variable $\rho$. For simplicity, we shall continue to work with the single field variable $\psi$, together with its corresponding momenta $\Pi^{\mu}_{\psi}$; the generalisation to multiple fields requires the trivial replacement $\psi\rightarrow\psi^{a}$ and $\Pi^{\mu}_{\psi}\rightarrow\Pi^{\mu}_{a}$, with $a=1,\,\cdots,k$.
Suppose that $\gamma:M\rightarrow J^{1}\widetilde{E}^{*}\times\mathbb{R}^{d}$ is a critical section of the multicontact variational problem, that we write in local coordinates as $\gamma(x)=\left(x^{\mu},\psi(x),\Pi^{\mu}_{\psi}(x),s^{\mu}(x)\right)$; from (39), we have
$0$ $=\frac{\partial\Pi^{\mu}_{\psi}}{\partial x^{\mu}}+\left(\frac{\partial H^{c}}{\partial\psi}+\Pi^{\mu}_{\psi}\frac{\partial H^{c}}{\partial s^{\mu}}\right)$
$=\frac{\partial}{\partial x^{\mu}}\left(e^{-\rho}p^{\mu}_{\psi}\right)+\left(e^{-\rho}\frac{\partial H}{\partial\psi}+e^{-\rho}p^{\mu}_{\psi}\frac{\partial H^{c}}{\partial s^{\mu}}\right)$
$=e^{-\rho}\left(-\rho_{\mu}p^{\mu}_{\psi}+\frac{\partial p^{\mu}_{\psi}}{\partial x^{\mu}}+\frac{\partial H}{\partial\psi}+p^{\mu}_{\psi}\frac{\partial H^{c}}{\partial s^{\mu}}\right)$
$=e^{-\rho}\left(\frac{\partial p^{\mu}_{\psi}}{\partial x^{\mu}}+\frac{\partial H}{\partial\psi}\right)$
In passing to the final line, we have used (48), noting that if $\rho_{\mu}=-\frac{\partial L^{H}}{\partial s^{\mu}}$ then $\rho_{\mu}=\frac{\partial H^{c}}{\partial s^{\mu}}$.
@@
It is thus clear that the Herglotz De-Donder Weyl equation for $\Pi^{\mu}_{\psi}$ correctly implies the corresponding multisymplectic expression for the momenta $p^{\mu}_{\psi}$. The proof for the $\psi$ equation of motion proceeds analogously, and is thus left as an exercise. The dynamical evolution of the action density, which is given by
$\frac{\partial s^{\mu}}{\partial x^{\mu}}-\left(\Pi^{\mu}_{\psi}\frac{\partial H^{c}}{\partial\Pi^{\mu}_{\psi}}+H^{c}\right)=0$
is precisely the statement that $\partial_{\mu}s^{\mu}=L^{H}$, expressed in canonically-conjugate variables.
From our construction, we have identified a symmetry of the covariant configuration space $E$, which was lifted to a vector field on the first jet bundle $J^{1}E$. Upon adopting coordinates along the scaling direction, and performing a Legendre transform, we observed that the corresponding Hamiltonian scaling symmetry $D$ included, in addition to the pure scaling $\partial_{\rho}$, further logarithmic derivative terms for each of the momentum directions. Geometrically, the reduction process appears to take place on the multiphase space, since we make use of the Hamiltonian function $H$ and the 2-forms $\omega^{\mu}_{h}$, both of which are defined on $J^{1}E^{*}$. However, the effect of the elimination of the scaling variable $\rho$ is to produce a covariant configuration space $\widetilde{E}=E/\mathbb{R}_{+}$ of dimension $\textrm{dim}\,E-1$. The reduced Hamiltonian $H^{c}$ and 1-forms $\eta^{\mu}$ are not defined on the first jet bundle of this new configuration space, but rather a manifold whose topology is of the form $J^{1}\widetilde{E}^{*}\times\mathbb{R}^{d}$.
This occurs precisely because the $d$ momenta $p^{\mu}_{\rho}$, which are no longer canonically conjugate to any field variable, assume the role of the action density, spanning a Reeb distribution [65], thereby producing a reduced space which is a multicontact manifold. Within the Lagrangian formalism, we observe a similar process; in this case, rather than the momenta $p^{\mu}_{\rho}$, it is the $d$ coordinate derivatives $\rho_{\mu}$ which span the Reeb distribution, yielding a manifold of the form $J^{1}\widetilde{E}\times\mathbb{R}^{d}$.
V Field Theory Classification
In order to better understand the geometric insight offered by our symmetry reduction procedure, it is instructive to consider the different multisymplectic bundle structures that arise within a variety of physically interesting field theories. Since the vast majority of such models are gauge theories [86, 87], we shall attempt to construct a broad system of classification, based upon the gauge group, and thence deduce important features common to each type. Our discussion will closely follow that presented in [88] and [89]; the notation employed throughout remains consistent with that developed over the course of sections (II.2) and (II.3).
In general, we shall consider two principal classes of theory, whose distinction is determined by the relationship between the gauge group $G\subset\textrm{Aut}(E)$, and the spacetime diffeomorphism group $\textrm{Diff}(M)$. Given a diffeomorphism $f:M\rightarrow M$, we say that an automorphism $\Phi:E\rightarrow E$ covers $f$ if $\pi\circ\Phi=f\circ\pi$. Intuitively, the automorphism can be considered to ‘lift’ the diffeomorphism from the spacetime manifold to the configuration space, in a manner which is consistent with the action of $f$ on $M$ [90].
A theory is described as being parameterised if all (or a suitably large subset of) spacetime diffeomorphisms are covered by the transformations in $G$. Physically, this means that the theory is invariant under arbitrary coordinate reparameterisations; this is the case of General Relativity, for example, in which the entirety of $\textrm{Diff}(M)$ is covered by $G$.
A background theory, by contrast, refers to any system in which the gauge group is a subset of those automorphisms which cover only the identity on $M$. A Yang-Mills theory defined on a flat background spacetime, for instance, has as its gauge group a set of transformations which are completely internal, and thus unrelated to spacetime diffeomorphisms.

### v4

@@ V Contact Reduction of Lagrangian Systems
Here, the first equation, defining the action density, is inverted to eliminate reference to $\rho_{\mu}$ in the function $f$ from which $L^{H}$ is constructed. We would now like to verify that the Herglotz-Lagrange equation
$\frac{\partial}{\partial x^{\mu}}\frac{\partial L^{H}}{\partial\psi^{a}_{\mu}}-\frac{\partial L^{H}}{\partial\psi^{a}}-\frac{\partial L^{H}}{\partial\psi^{a}_{\mu}}\frac{\partial L^{H}}{\partial s^{\mu}}=0$ (34)
faithfully reproduces the dynamics of the original system, for the unscaled fields $\psi^{a}$. For this, we must make careful use of the chain rule, accounting for the dependence of $s^{\mu}$ on the other coordinates. We then have
$\begin{split}\frac{\partial L^{H}}{\partial\psi^{a}_{\mu}}=\frac{\partial f}{\partial\psi^{a}_{\mu}}+\frac{\partial f}{\partial\rho_{\nu}}\frac{\partial\rho_{\nu}}{\partial\psi^{a}_{\mu}}-s^{\nu}\frac{\partial\rho_{\nu}}{\partial\psi^{a}_{\mu}}\\ \frac{\partial L^{H}}{\partial\psi^{a}}=\frac{\partial f}{\partial\psi^{a}}+\frac{\partial f}{\partial\rho_{\nu}}\frac{\partial\rho_{\nu}}{\partial\psi^{a}}-s^{\nu}\frac{\partial\rho_{\nu}}{\partial\psi^{a}}\\ \frac{\partial L^{H}}{\partial s^{\mu}}=\frac{\partial f}{\partial\rho_{\nu}}\frac{\partial\rho_{\nu}}{\partial s^{\mu}}-\rho_{\mu}-s^{\nu}\frac{\partial\rho_{\nu}}{\partial s^{\mu}}\end{split}$
@@
Combining these results, recalling that $s^{\mu}=\partial f/\partial\rho_{\mu}$, the left-hand side of (34) becomes
$\frac{\partial}{\partial x^{\mu}}\frac{\partial f}{\partial\psi^{a}_{\mu}}-\frac{\partial f}{\partial\psi^{a}}+\rho_{\mu}\frac{\partial f}{\partial\psi^{a}_{\mu}}$
Multiplying through by $e^{\rho}$, we recover precisely the multisymplectic field equation for $\psi^{a}$
$\frac{\partial}{\partial x^{\mu}}\left(e^{\rho}\frac{\partial f}{\partial\psi^{a}_{\mu}}\right)-e^{\rho}\frac{\partial f}{\partial\psi^{a}}=0$
We have thus shown that the Herglotz Lagrangian (33) correctly reproduces the dynamics of the original system, whilst making no reference to $\rho$, confirming that this scaling variable plays no role in the dynamical evolution of the observable degrees of freedom.
VI Contact Reduction of Hamiltonian Systems
Suppose that $(J^{1}E^{*},\Omega_{H})$ is a multisymplectic Hamiltonian system, with corresponding Hamiltonian function $H:J^{1}E^{*}\rightarrow\mathbb{R}$. For simplicity, we shall assume that the Legendre map is a global diffeomorphism, so that our system is hyperregular. A vector field $D\in\mathfrak{X}^{\infty}(J^{1}E^{*})$ is said to constitute a scaling symmetry of degree $\Gamma\in\mathbb{R}$ if
$\mathfrak{L}_{D}H=\Gamma H\quad\quad\quad\textrm{and}\quad\quad\quad\mathfrak{L}_{D}\hskip 0.85358pt\omega^{\mu}_{H}=\omega_{H}^{\mu}\quad\quad\textrm{for all}\;\mu=0,\cdots\,,d-1$
in which the 2-forms $\omega_{H}^{\mu}$ were defined in (9).
@@ VI Contact Reduction of Hamiltonian Systems
The Hamiltonian reduction may, in principle, be carried out without reference to a Lagrangian function. However, in practice, it is more convenient to work with the Hamiltonian obtained via the Legendre transform of $L=e^{\rho}f(\rho_{\mu},\psi^{a},\psi^{a}_{\mu})$, for this implies that
$\mathcal{FL}^{*}p^{\mu}_{\rho}=\frac{\partial L}{\partial\rho_{\mu}}=e^{\rho}\frac{\partial f}{\partial\rho_{\mu}}=e^{\rho}s^{\mu}\quad\quad\quad\quad\mathcal{FL}^{*}p^{\mu}_{a}=\frac{\partial L}{\partial\psi^{a}_{\mu}}=e^{\rho}\frac{\partial f}{\partial\psi^{a}_{\mu}}$ (35)
@@
By virtue of the structure of the Lagrangian scaling symmetry $\Sigma=\partial_{\rho}$, we find that the corresponding vector field $D$ for the Hamiltonian is
$D=\frac{\partial}{\partial\rho}+p^{\mu}_{\rho}\frac{\partial}{\partial p^{\mu}_{\rho}}+p^{\mu}_{a}\frac{\partial}{\partial p^{\mu}_{a}}$ (36)
Here, the De-Donder Weyl formalism greatly facilitates our construction. The scaling symmetry must act on both the temporal and spatial momenta; the latter do not feature amongst the phase space variables in the canonical framework. On the reduced space (the details of which will be discussed shortly) we introduce coordinates
$s^{\mu}=\frac{p^{\mu}_{\rho}}{e^{\rho}}\hskip 93.89418pt\Pi_{a}^{\mu}=\frac{p^{\mu}_{a}}{e^{\rho}}$ (37)
in which, for simplicity, we have omitted the pullback via the projection map. Finally, we claim that, upon defining
$H^{c}(\psi^{a},\Pi_{a}^{\mu},s^{\mu})=\frac{H}{e^{\rho}}\hskip 93.89418pt\eta^{\mu}(\psi^{a},\Pi^{\nu}_{a},s^{\nu})=\frac{\iota_{D}\hskip 0.56905pt\omega^{\mu}_{H}}{e^{\rho}}=ds^{\mu}-\Pi^{\mu}_{a}\,d\psi^{a}$ (38)
@@
we obtain a multicontact structure, whose corresponding Hamiltonian $d$-form $\Theta_{H^{c}}$ is given by
@@
Further, the contact Hamiltonian equations (23) for holonomic sections correctly reproduce the dynamics of the full multisymplectic system, without reference to the scaling variable $\rho$. In order to prove our claim, we shall suppose that $\gamma$ is a holonomic section, written locally as $\gamma(x)=(x^{\mu},\psi^{a}(x),\Pi^{\mu}_{a}(x),s^{\mu}(x))$; from (23), the equation of motion for the momenta $\Pi^{\mu}_{a}$ reads
$0$ $=\frac{\partial\Pi^{\mu}_{a}}{\partial x^{\mu}}+\left(\frac{\partial H^{c}}{\partial\psi^{a}}+\Pi^{\mu}_{a}\frac{\partial H^{c}}{\partial s^{\mu}}\right)$
$=\frac{\partial}{\partial x^{\mu}}\left(e^{-\rho}p^{\mu}_{a}\right)+\left(e^{-\rho}\frac{\partial H}{\partial\psi^{a}}+e^{-\rho}p^{\mu}_{a}\frac{\partial H^{c}}{\partial s^{\mu}}\right)$
$=e^{-\rho}\left(-\rho_{\mu}p^{\mu}_{a}+\frac{\partial p^{\mu}_{a}}{\partial x^{\mu}}+\frac{\partial H}{\partial\psi^{a}}+p^{\mu}_{a}\frac{\partial H^{c}}{\partial s^{\mu}}\right)$
$=e^{-\rho}\left(\frac{\partial p^{\mu}_{a}}{\partial x^{\mu}}+\frac{\partial H}{\partial\psi^{a}}\right)$
@@
In passing to the final line, we have used (30), noting that if $\rho_{\mu}=-\frac{\partial L^{H}}{\partial s^{\mu}}$ then $\rho_{\mu}=\frac{\partial H^{c}}{\partial s^{\mu}}$. It is thus clear that the contact Hamiltonian equation for $\Pi^{\mu}_{a}$ implies the corresponding multisymplectic expression for the momenta $p^{\mu}_{a}$. The proof for the $\psi^{a}$ equation of motion proceeds analogously, and is thus left as an exercise.
Having presented both the Lagrangian and Hamiltonian constructions of the symmetry-reduced theory, we close with a discussion of the geometry of the multicontact spaces upon which $L^{H}$ and $H^{c}$ are defined. Within the original multisymplectic Lagrangian system, we identified a symmetry of the covariant configuration space $E$, which was lifted to a vector field on the first jet bundle $J^{1}E$. Upon adopting coordinates along the scaling direction, we implicitly revealed that $E$ could be written as the sum of two connected pieces $E_{\pm}=(M\times\mathbb{R}_{\pm})\times_{M}\widetilde{E}$, in which $\widetilde{E}$ is a codimension-one subspace of $E$, identified as a configuration space comprising all unscaled field variables. $M\times\mathbb{R}_{\pm}\rightarrow M$ is a trivial bundle over $M$, and relates to a subtlety not yet discussed. Upon making a change of coordinates, the scaling vector field adopted the form (31); under the subsequent identification $\xi=e^{\rho/\Lambda}$, we implicitly assumed that $\xi>0$, which may not be the case. As such, in order that no dynamical information be lost, when the scaling symmetry has been rendered of the form (31), we should consider both $\xi=e^{\rho/\Lambda}$ and $\xi=-\,e^{\rho/\Lambda}$. It is precisely these choices which yield the trivial bundle $M\times\mathbb{R}_{\pm}\rightarrow M$.
Provided both components $M\times\mathbb{R}_{\pm}$ are considered, it is clear that the quotient space under orbits of $\Sigma$ is simply $\widetilde{E}$. The reduction process eliminates only the scaling variable $\rho$, and not its associated velocities $\rho_{\mu}$; recall that it was necessary to use the expression (33) to eliminate reference to $\rho_{\mu}$ in favour of the action density. As a result, the reduction takes place at the level of the configuration space, and so the Herglotz Lagrangian is defined on $J^{1}\widetilde{E}\times\mathbb{R}^{d}$, in which the additional factor of $\mathbb{R}^{d}$ corresponds to the components of the action density, which have replaced the $\rho_{\mu}$. Within the Hamiltonian formalism, the idea is similar; however, it is now the momenta $p_{\rho}^{\mu}$ that assume the role of the action density, and so the reduced space upon which the contact Hamiltonian is defined is of the form $J^{1}\widetilde{E}^{*}\times\mathbb{R}^{d}$.
VII Field Theory Classification
The vast majority of physically interesting models are gauge theories [44, 45]; at present, our formalism is suited only to the reduction of regular theories. However, in anticipation of a need to have at our disposal a system of classification, based upon the gauge group, we shall briefly discuss the various bundle structures that arise in different kinds of theory. Our presentation follows closely that of [46] and [47].